Calculate the milligrams of arabinose required to get a 5% solution dissolved in 10 mL water. (Arabinose MW: 150.13 g/mol).
Answer #1

5% arabinose means 5 gram arabinose is dissolved in 100 ml solution.

We need only 10 ml of 5% solution.

So,

100 ml = 5 gram

10 ml = 5/100 × 10 = 0.5 gram arabinose

0.5 gram = 0.5 × 1000 = 500 mg

We need 500 mg arabinose to prepare 5%, 10 ml of its solution.

Please rate.